Actions
20170530 » History » Revision 2
« Previous |
Revision 2/4
(diff)
| Next »
Rajaram, Durga, 30 May 2017 13:10
Ops/Comm Scheduling meeting : 30/5/2017¶
Location : CR21 at 14:00 pm
Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 1779339517
Agenda¶
- Actions
- Macwaters - Remove old overheated CSW71 to R9 waiting for John Govans and the van - in progress
- Macwaters - Report on status of FC psu repair (stands)
- The spare is installed and has been run into a short
- Hall Manager - water system change-over procedure (stands) - collar people on Friday
- EOverton - Status of
miceraid5
disk cleanup (stands) now has a script but needs to run it - Progress on Hall list List of Hall activities in the shutdown (Colin) DL were down on masse and made significant progress. Need to confirm if any remaining items
- CWhyte to request survey of TOF1 - requested, due Thursday Removing absorber means will need to re-survey CC magnets after the User Cycle.
- AKurup to check which settings are being used by the active Alarm Handlers
- CWhyte to request survey of CC Magnets after cycle
- FDrielsma to investigate low-efficiency channels in EMR x-z
- Tracker group: decide what should be done to prevent repeat contamination of Tracker Cryostats
- MOM update
- Performance plots from TOF calibration run (Durga?) Run9376imageData.pdf
- Hydrogen system update
- Neon test
- Run plan
- AOB
- Last meeting : Tuesday May 9 2017
- Next meeting : Tuesday June 6 2017
Updated by Rajaram, Durga over 6 years ago · 2 revisions